# Lemmanaid: Neuro-Symbolic Lemma Conjecturing

### v3

@@ A.6 Partial QuickSpec output on Octonions
QuickSpec takes as input a list of function and constant symbols out of which it will build the terms. We started by giving it the functions (with $\times$, inverse and $1$). As the inverse of 0 is not defined, we had to instruct QuickSpec to only test using non-zero Octonions, a limitation not shared by Lemmanaid. For the same reason we could not give it the constant symbol 00 or the functions $+$ and --, which would allow it to construct a zero octonion. The output of QuickSpec was as follows:
@@
Then we ran QuickSpec again, adding the extra functions $+$, 00, inner product ($\cdot$) and norm, but removing the inverse operation (as discussed above). The results are still reasonable but, especially with the inner product function, most of the lemmas are uninteresting, not found in the AFP theory, and generated only because they happen to be true:
@@
[…]
A.7 Example Lemmas Discovered by Lemmanaid
There are numerous examples of Lemmanaid predicting correct templates (and then recovering the gold-standard lemmas), but it is sometimes difficult to explain exactly why Lemmanaid succeeded when other methods failed. In Octonions, for example, Lemmanaid recovers the following lemma, which neither the neural baseline nor QuickSpec recovers:
    lemma Im7_tendsto_lowerbound:
   "\<lbrakk>
      (f \<longlongrightarrow> limit) net;
      \<forall>\<^sub>F x in net. b \<le> Im7 (f x);
      net \<noteq> bot
    \<rbrakk>
      \<Longrightarrow> b \<le> Im7 limit"
The input (below) associated with this lemma is truncated when sent to the LLM due to our max sequence length, but Lemmanaid is still able to recover the correct template.
###symbols
  Orderings.bot_class.bot
  Octonions.octo.Im7
  Filter.eventually
  Topological_Spaces.topological_space_class.tendsto
###defs
class bot =
  fixes bot :: ’a ("\<bottom>")
codatatype octo =
  Octo (Ree: real) (Im1: real) (Im2: real) (Im3: real) (Im4: real)
       (Im5: real) (Im6: real) (Im7: real)
definition eventually :: "(’a \<Rightarrow> bool) \<Rightarrow> ’a filter
  \<Rightarrow> bool"
  where "eventually P F \<longleftrightarrow> Rep_filter F P"
class topological_space = "open" +
  assumes open_UNIV [simp, intro]: "open UNIV"
  assumes open_Int [intro]: "open S \<Longrightarrow> open T \<Longrightarrow>
        open (S \<inter> T)"
  assumes open_Union [intro]: "\<forall>S\<in>K. open S \<Longrightarrow>
        open (\<Union>K)"
Lemmanaid’s neural engine predicts the following (correct) template:
    \<lbrakk>
      ?H1 x_1 x_2 x_3;
      ?H2 (\<lambda>y_0. x_4 \<le> ?H3 (x_1 y_0)) x_3;
      x_3 \<noteq> ?H4
    \<rbrakk>
      \<Longrightarrow> x_4 \<le> ?H3 x_2
Templates are typically shorter than their corresponding lemma objects, which may help with better reasoning in some cases for neural models. The following is an example lemma object from HOL-test that our neural baseline could not predict directly:
    \<lbrakk>
      continuous_map ?X ?Y ?f; compact_space ?X;  Hausdorff_space ?Y ;
      ?f ‘ topspace ?X = topspace ?Y;  inj_on ?f (topspace ?X)
    \<rbrakk>
      \<Longrightarrow> homeomorphic_map ?X ?Y ?f
The above lemma object is 230 characters, but the template extracted from this lemma is only 134. The associated template, which Lemmanaid’s neural engine is able to correctly predict and then its symbol engine is able to successfully instantiate, looks as follows:
    \<lbrakk>
      ?H1 x_1 x_2 x_3; ?H2 x_1; ?H3 x_2;
      ?H4 x_3 (?H5 x_1) = ?H6 x_2; ?H7 x_3 (?H5 x_1)
    \<rbrakk>
      \<Longrightarrow> ?H8 x_1 x_2 x_3
A.8 Successes and failures of Lemmanaid
We provide a few examples of conjecturing problems attempted by Lemmanaid, covering some successful and failed attempts.
Success
Let’s consider the lemma BExp.bval_not which features in IMP/BExp theory in HOL:
    lemma bval_not[simp]: "bval (not b) s = (\<not> bval b s)"
    |gold-standard| ?H1 (?H2 x_1) x_2 = (\<not> ?H1 x_1 x_2)
    |prediction| ?H1 (?H2 x_1) x_2 = (\<not> ?H1 x_1 x_2)
Lemmanaid predicts an identical template to the gold-standard, and the template can be successfully instantiated to generate a conjecture which is equivalent to bval_not.
Success with Many Symbols
Lemmanaid is able to handle conjecturing problems with a lot of symbols, which can be tricky for symbolic tools which exhaustively enumerate conjectures. For example, consider the lemma logderiv_zeta_region_estimate from an AFP formalization of prime number theory with remainder term:
    lemma logderiv_zeta_region_estimate:
      assumes "s \<in> logderiv_zeta_region"
      shows "\<parallel>logderiv zeta s\<parallel> \<le> C\<^sub>2 *
            (ln (\<bar>Im s\<bar> + 3))\<^sup>2"
    |gold-standard| x_1 \<in> ?H1 \<Longrightarrow>
    ?H2 (?H3 ?H4 x_1) \<le> ?H5 ?H6 (?H7 (?H8 (?H9 (?H10 (?H11 x_1))
    (?H12 (?H13 ?H14)))))
    |prediction| x_1 \\<in> ?H1 \\<Longrightarrow>
    ?H2 (?H3 ?H4 x_1) \\<le> ?H5 ?H6 (?H7 (?H8 (?H9 (?H10 (?H11 x_1))
    (?H12 (?H13 ?H14)))))
Generating the correct template here requires successfully relating 14 symbols, and reasoning over their corresponding types and definitions.
Off-target (But Valid) Prediction
Lemmanaid sometimes generates helpful templates that differ from gold-standard templates. Such templates are still useful, as they can be instantiated into valid lemmas, but our evaluation considers such templates as misses. Consider attempting to conjecture the lemma bval_and:
    lemma bval_and[simp]: "bval (and b1 b2) s = (bval b1 s \<and> bval b2 s)"
    |gold-standard| ?H1 (?H2 x_1 x_2) x_3 = (?H1 x_1 x_3 \<and> ?H1 x_2 x_3)
    |prediction| \<lbrakk> ?H1 x_1 x_2; ?H1 x_3 x_2 \<rbrakk> \<Longrightarrow>
    ?H1 (?H2 x_1 x_3) x_2
In this case, the predicted template is instantiated to a different, valid lemma, namely \<lbrakk> bval b1 s; bval b2 s \<rbrakk> \<Longrightarrow> bval (and b1 b2) s. This conjectured lemma is weaker lemma than the gold-standard, but it is equivalent to assuming the right-hand side of the equality to prove the left-hand side. Counter-example checking and other property testing techniques could be leveraged in these cases to filter valid candidates.
Invalid Prediction
Consider the totient function, which counts the number of positive integers up to an input n which are relatively prime to n. One property of the totient is:
    lemma totient_le: "totient n \<le> n"
    |gold-standard| ?H1 x_1 \<le> x_1
    |prediction| x_1 \\<le> ?H1 x_1
In this case, Lemmanaid predicted a similar template to the gold-standard template, though the lemma created from instantiating this template is invalid. Again, counterexample checking can help filter invalid lemma candidates.

### v5

@@ Appendix H Partial QuickSpec output on Octonions
QuickSpec takes as input a list of function and constant symbols out of which it will build the terms. We started by giving it the functions (with $\times$, inverse and $1$). As the inverse of 0 is not defined, we had to instruct QuickSpec to only test using non-zero Octonions, a limitation not shared by Lemmanaid. For the same reason we could not give it the constant symbol 0 or the functions ++ and -, which would allow it to construct a zero octonion. The output of QuickSpec is shown in Figure 2.
Figure 2: Output from QuickSpec on Octonions for multiplication and inverse.
@@
Then we ran QuickSpec again, adding the extra functions ++, 0, inner product ($\cdot$) and norm, but removing the inverse operation (as discussed above). The results are still reasonable but, especially with the inner product function, most of the lemmas are uninteresting, not found in the AFP theory, and generated only because they happen to be true (see Figures 3 and 4).
Figure 3: Output from QuickSpec on Octonions with ++, 0, inner product ($\cdot$)
@@
Figure 4: Continued output from Figure 3, Octonions with ++, 0, inner product ($\cdot$) and norm
@@
In the complex numbers, the functions $\operatorname{Re},\operatorname{Im}:\mathbb{C}\rightarrow\mathbb{R}$ extract the real and imaginary parts of a number respectively. In the octonions, there are eight analogous functions $\operatorname{Re},\operatorname{Im_{1}}\ldots\operatorname{Im_{7}}:\mathbb{O}\rightarrow\mathbb{R}$. When we added these functions, QuickSpec generated 775 laws, the vast majority extremely uninteresting for a human. An example of a typical generated law is $\operatorname{Re}(x+\operatorname{Im_{2}}\,y\times z)=\operatorname{Re}(x+y\times\operatorname{Im_{2}}\,z)$.
Corresponding to the complex number $i$ there are seven octonions $e_{1}\ldots e_{7}$ which (together with $1$) act as unit numbers. Unfortunately when we add these as constants then QuickSpec starts to generate many thousands of irrelevant lemmas. Examples include:
$(e_{5}-e_{4})\times(e_{1}+e_{4})=(1+e_{1})\times(e_{4}-e_{1})$
and
$e_{4}\times(\operatorname{Im_{1}}\,x\times\operatorname{Im_{2}}\,y)=\operatorname{Im_{7}}(\operatorname{Im_{1}}\,x\times(y\times e_{4}))$
The problem is that QuickSpec does not attempt to judge which lemmas are relevant to a human user.